OpenStudy (anonymous):

simplify: (15mn^2p)/25m^2n^2

OpenStudy (gw2011):

If you mean: (15mn^2)(p)/25m^2n^2, then you can do the following: 15/25=3/5 m/m^2=1/m n^2/n^2=1 From this you get: 3p/5m